A uniform plane wave travelling in air is incident on the plane boundary between air and another dielectric medium with εr = 4. The reflection coefficient for the normal incidence is


1
Expert's answer
2021-01-07T10:56:17-0500

In general we have

"n = \sqrt{\epsilon_r \mu_r}"

For dielectric medium we can suppose "\mu_r \approx 1" and thus "n =\sqrt{\epsilon_r}" .

Therefore we have "\rho = \frac{(n-1)^2}{(n+1)^2}=\frac{1}{9}" .